I am using an online study guide for the COMPASS Assessment Test, I am having trouble with this particuluar word problem. The problem involves d= rt.
If 2 planes leave the same airport at 1:00 PM, how many miles apart will they be at 3:00 PM, if one travels directly north at 150 mph and the other plane travels directly west at 200 mph.
A. 50 miles
B. 100 miles
C. 500 miles-correct answer
D. 700 miles
E. 1000 miles
I originally answered the question D 700 miles, by taking 150+ 200, multiplying the answer by 2 for 2 hours.

Found 2 solutions by stanbon, scott8148:
Answer by stanbon(48551)

If 2 planes leave the same airport at 1:00 PM, how many miles apart will they be at 3:00 PM, if one travels directly north at 150 mph and the other plane travels directly west at 200 mph.
-------------------------
North traveler goes 2*150 = 300 miles
West traveler goes 2*200 = 400 miles
-----------------------------------------
Use Pytagoras to find the distance of one plane to the other.
d = sqrt[300^2 + 400^2] = sqrt[250000] = 500 miles
===========================
Cheers,
Stan H.
============================

A. 50 miles
B. 100 miles
C. 500 miles-correct answer
D. 700 miles
E. 1000 miles

Answer by scott8148(5890)

the two planes are flying at a right angle to each other (north and west)

their distances from the airport are the legs of a right triangle, and the distance from each other is the hypotenuse

north bound __ d=rt __ d=150*2 __ d=300

west bound __ d=rt __ d=200*2 __ d=400

so the distance between them (hypotenuse) is ... think 3-4-5 __ or a^2+b^2=c^2